# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### v1

@@ 5.3. Reinforcement Learning Methods
Reinforcement learning (RL), a subfield of machine learning, is a powerful approach that enables an agent to learn how to make optimal decisions automatically (Sutton and Barto, 2018). It has been successfully applied in various domains, including robotics, autonomous vehicles, education, and healthcare (Arulkumaran et al., 2017; Kober et al., 2013). In RL, an agent interacts with an environment and receives feedback in the form of rewards or penalties based on its actions. The goal is to learn a policy $\pi$, which can maximize the long-term cumulative reward. As shown in Figure 6, the policy can be learned by exploring the environment and learning from its consequences of actions. In essence, researchers in CAT utilize RL methodologies to address a question: Can the selection algorithm (policy) be automatically learned and optimized from data or student interactions, thus circumventing the necessity for expert intervention?
Markov Decision Process Formulation. The interaction between agent and environment can be viewed as a Markov Decision Process (MDP) (Feinberg and Shwartz, 2012). Specifically, at each step, the agent observes current environment’s state ($s$), and interacts with the environment by selecting its actions ($a$). Simultaneously, the agent receives a reward ($r$) from these interactions, influencing or changing the current state of the environment. The objective is to select a best sequence of actions, resulting in the highest cumulative reward ($\sum r$). Therefore, most RL problems are formally described as estimating the optimality of the agent’s behavior in a given state (value-based methods (Sutton and Barto, 2018)) or the optimality of the action policy itself (policy-based methods (Agarwal et al., 2021)) or the hybrid approaches (Mnih et al., 2016). We next formulate the CAT problem as an MDP. Here are the definitions of the key elements in RL within an adaptive learning/testing system and the overall RL framework of CAT is in Figure 6:
@@
For modeling the complex interactions between examinees and questions during the practical test, Zhuang et al. (2022b) recently design a transformer-based Q-Network, which can identify and extract the perturbations in examinee performance (e.g., guess and slip factors).
Stochastic Shortest Path Formulation. Furthermore, CAT can be defined as a Stochastic Shortest Path (SSP) problem (Bertsekas and Tsitsiklis, 1991), which is a special case of MDP. In an SSP, the objective is to find the shortest path (i.e., the minimum test step) from a given initial state $s_{0}$ to goal states. In CAT, the goal state typically represents the completion of the test or the attainment of a predetermined level of proficiency estimation precision. The approach by Gilavert and Freire (2022) uses Linear Programming to find the optimal testing policy $\pi^{*}$, treating the problem like a flow network where each state must have balanced inflow and outflow, except for the start and end points. It denotes variables $x_{s,a}$ as the expected accumulated occurrence frequency for every pair (state $s\in S$, action $q\in\mathcal{Q}$), and equalizes $in(s)$ and $out(s)$ flow model for every state $s$. The flow into a state $s$ is the sum of the expected frequencies of all actions in all other states $s^{\prime}$ that lead to $s$: ${in}(s)=\sum_{s^{\prime},q}x_{s^{\prime},q}P(s|s^{\prime},q)$. The flow out of a state $s$ is the sum of the expected frequencies of all actions in state $s$: ${out}(s)=\sum_{q}x_{s,q}$. The objective function is to maximize the total expected reward $r$, which is the sum of the expected frequencies times the immediate rewards $r(s,q)$ for all state-action pairs: $\min_{x_{s,q}}\sum_{s\in S,q\in\mathcal{Q}}{x_{s,q}r(s,q)}$. Thus the optimal policy $\pi^{*}$ can be obtained by:
@@
Partial-Observable MDP Formulation. Partial-Observable MDP (POMDP) is a generalization of MDP to handle situations where the agent does not fully observe the environment (Hoerger and Kurniawati, 2021). Previously, it assumes that the proficiency of an examinee can be perfectly inferred from their previous responses, then CAT can be modeled as an MDP, where the proficiency estimate can be the states. However, in a more general scenario, the proficiency cannot be perfectly inferred from their previous responses. Because there is always some uncertainty involved (Zhuang et al., 2022a):
@@ 5.4. Meta Learning Algorithms
Figure 7(left) shows the overall meta learning framework. In the inner-level (Eq.(12)), the question in the support set $\mathcal{D}_{s}^{i}$ for examinee $i$ is sequentially selected by $\pi$, according to his/her previous responses; then binary cross-entropy loss $l(\cdot)$ on $\mathcal{D}_{s}^{i}$ is minimized for estimating the proficiency $\hat{\theta}_{i}$ for the outer-level. In the outer-level (Eq.(11)), the loss of the estimate $\hat{\theta}_{i}$ on the query set $\mathcal{D}_{u}^{i}$ is minimized to learn the selection algorithm $\pi$ and the global parameters $\gamma$. The global parameters $\gamma$ refer to the question characteristics and the examinee ability prior. The algorithm $\pi$ is also model-agnostic. It could be adapted to the given CDM ($f$) automatically by optimizing this problem for efficient selection. Once the question selection algorithm is trained, its parameters do not update during the CAT process and adaptively selects the next item based on previous response behaviors.
Based on the above bi-level definition, there have been increasing efforts to improve upon it. Ma et al. (2023) propose a flexible optimization framework Decoupled Learning CAT (DL-CAT). The original BOBCAT obtains the parameters of two modules (i.e., examinee proficiency estimation and question selection algorithm) through coupled inner and outer optimizations, i.e., the result of the outer optimization model is used to measure the quality of the inner. DL-CAT devises a ground-truth construction strategy, and a pairwise loss function, allowing these two models to be trained independently; Feng et al. (2023) introduces a constrained version of BOBCAT to address the question exposure and test overlap issues. Yu et al. (2023) recently introduce the collaborative information of examinees in optimizing this bi-level problem, achieving fast convergence of proficiency estimation.
Meta-Learning vs Reinforcement Learning. In CAT, meta-learning can be seen as a higher-level learning process that learns how to adapt a general strategy for question selection to specific examinees based on their responses. Actually, this meta-learning CAT can be reframed as a RL problem. Zhuang et al. (2022b) propose NCAT which transforms the meta-learning problem in CAT into a RL problem. Because the test may stop at any step according to different stopping rules, NCAT simplifies the original objective (Eq(11)) in meta-learning and sums all the test steps to minimize the loss:
@@ 5.5. Subset Selection Algorithms
where $\hat{\theta}^{T}=\arg\min_{\theta}{\sum_{(q,y)\in S}{l(y,f(q,\theta))}}$ is the final proficiency estimate when the test ends $(t=T)$ with the corresponding $T$ responses. In contrast to previous sequential selection methods, it essentially does not require perfect selection at each step, but rather emphasizes the accuracy of the final estimate.
From a global perspective, CAT essentially is a Subset Selection problem (Miller, 2002), a fundamental challenge in machine learning and optimization. It revolves around choosing a subset of elements $S$ from a larger set $\mathcal{Q}$ that optimizes a particular objective function $F(S)$ while adhering to specific constraints. However, we cannot directly solve the above optimization problem due to the following main challenge: The true proficiency of the examinee, denoted by $\theta_{0}$, is unknown. It is not available in the dataset, which prevents us from directly optimizing or designing the question selection algorithm. To address this, some researchers have tried to approximate the original objective. Mujtaba and Mahapatra (2021) use the standard error of measurement (SEM) to approach the optimization objective. The SEM provides a measure of confidence in an estimate from a test. They propose using SEM as a new target to measure the accuracy of proficiency estimation: $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}},$ where $\mathcal{I}_{i}(\hat{\theta})$ is the information (e.g., Fisher Information) for a question $i$ for an examinee with proficiency $\hat{\theta}$, thus the objective function $F(S)=\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}}$. At each step, it uses multi-objective evolutionary algorithms to obtain the set of Pareto-optimal solutions (Deb, 2011) by maximizing precision and minimizing the number of questions.
@@
where $w(i,j)\triangleq d-\mathrm{max}_{\theta\in\Theta}{\|\nabla l_{i}(\theta)-\nabla l_{j}(\theta)\|}$ is the gradient similarity between question pair $(q_{i},q_{j})$ for this examinee, thus the objective function $F(S)=\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j)$. The core of BECAT’s subset selection strategy is to find a subset $S$ of size $T$ that maximizes the coverage of $\mathcal{Q}$, quantified by the similarity measure $w(i,j)$ between each question $i\in\mathcal{Q}$ and its closest counterpart $j\in S$. This approach, represented in Figure 8, essentially seeks the most representative questions, aligning with prior selection algorithms but under a new, more rigorous theoretical framework.
Given the NP-Hard nature of this optimization, BECAT employs a submodular function approximation. A simple greedy algorithm can finally solve this subset selection problem, with BECAT ensuring that the estimate error remains upper-bounded at each step. The subset selection problem in CAT is a fresh direction with significant potential. This method offers a universal and global framework for question selection, applicable across various complex CDMs that can utilize gradient-based estimations, including neural network models. Crucially, it also provides a theoretical basis for estimation error analysis.
@@
Statistical    Algorithms Simple implementation and efficient operation Dependent on IRTs and requires expert knowledge for design
Active Learning Model-agnostic and flexible Neglect the nuanced information within CDM parameters
Reinforcement Learning Automatic generation of selection algorithm; Sequential Decision Making Incurs additional training costs and potential bias from data-driven selection
Meta Learning Automatic generation of selection algorithm; Fast Adaptation Incurs additional training costs and potential bias from data-driven selection
Subset Selection Strong theoretical guarantees for estimation accuracy Faces challenges in the initial stages of CAT
@@
Table 1 compares these five categories of question selection algorithms in CAT systems, highlighting the generality and interpretability of each category, along with their main advantages and limitations. This overview assists researchers in identifying the most suitable algorithm for their CAT applications, balancing efficiency and complexity (a summary of the performance of different methods is presented in Table 3).
@@ 7.2. Fairness
Fairness is a topic of profound societal significance in both education and machine learning research fields, sparking numerous discussions and leading to the development of many fairness-aware learning algorithms (Cleary, 1968; Chai and Wang, 2022; Li and Liu, 2022; Jiang and Nachum, 2019). As a technology with potential applications in high-stakes testing, fairness in CAT is a paramount concern. In CAT, the bias that leads to fairness issues can be introduced through three components: the Cognitive Diagnosis Model (CDM), the selection algorithm, and the question bank:
@@ 7.4. Search Efficiency
In large-scale educational testing, such as the GRE (Mills and Steffen, 2000), the search efficiency of selection algorithms poses a significant challenge. The algorithms typically evaluate the value of each question in a brute-force manner. Consequently, the time complexity of selecting a question in each testing step is linear to the number of questions in the bank $\mathcal{Q}$, denoted as $O(|\mathcal{Q}|)$. This linear relationship is a bottleneck in intelligent educational systems. Many organizations like GMAT (Rudner, 2009) resort to the manual development of filtering rules by experts, a labor-intensive process that somewhat limits the efficacy of CAT. To address the efficiency issue in question selection, researchers have proposed two kinds of solutions: (1) The application of the Particle Swarm Optimization (PSO) algorithm in IRT has proven effective in mitigating computational complexities (Huang et al., 2009). PSO operates with particles concurrently exploring the search space, where each particle represents a potential solution to the optimization problem of question selection. This parallel exploration allows PSO to cover a larger portion of the search space in each iteration, improving the chances of rapidly identifying optimal questions; (2) Leveraging efficient search methods from the realms of recommendation systems and information retrieval to enhance CAT’s search efficiency, e.g., creating of a tree-based index (Bentley, 1975; Ram and Gray, 2012; Zhu et al., 2018). This approach organizes data in hierarchical tree structures, with balanced trees ensuring logarithmic time complexity for search operations. Recently, Hong et al. (2023) propose a Search-Efficient CAT framework, which employs examinee-aware space partitioning to construct a tree-based index. It reduces the search space and eliminates redundant searches in multi-round testing steps, thereby reducing the search complexity from $O(|\mathcal{Q}|)$ to $O(\log|\mathcal{Q}|)$.
@@
Content Control Exposure
@@
Unbalanced question usage Mitigates overexposure;
@@
Test Coverage Algorithmic preferences Comprehensive assessment;
@@
Fairness Bias in CDMs Skewed training data;
@@
Algorithmic preferences Reduces disadvantage for certain groups (Lord, 2012)
Equating Different selected questions across examinees Score comparability;
@@
Robustness Noise Resistance Random variability;
@@
Modeling Uncertainty Uncertainty in response Improves accuracy of proficiency estimation (Veldkamp and Verschoor, 2019)
@@
Search Efficiency Large question banks;

### v2

@@ 5.3. Reinforcement Learning Methods
Reinforcement learning (RL), a subfield of machine learning, is a powerful approach that enables an agent to learn how to make optimal decisions automatically (Sutton and Barto, 2018). It has been successfully applied in various domains, including robotics, autonomous vehicles, education, and healthcare (Arulkumaran et al., 2017; Kober et al., 2013). In RL, an agent interacts with an environment and receives feedback in the form of rewards or penalties based on its actions. The goal is to learn a policy $\pi$, which can maximize the long-term cumulative reward. As shown in Figure 6, the policy can be learned by exploring the environment and learning from its consequences of actions. In essence, researchers in CAT utilize RL methodologies to address a question: Can the selection algorithm (policy) be automatically learned and optimized from data or student interactions, thus circumventing the necessity for expert intervention? Markov Decision Process Formulation. The interaction between agent and environment can be viewed as a Markov Decision Process (MDP) (Feinberg and Shwartz, 2012). Specifically, at each step, the agent observes current environment’s state ($s$), and interacts with the environment by selecting its actions ($a$). Simultaneously, the agent receives a reward ($r$) from these interactions, influencing or changing the current state of the environment. The objective is to select a best sequence of actions, resulting in the highest cumulative reward ($\sum r$). Therefore, most RL problems are formally described as estimating the optimality of the agent’s behavior in a given state (value-based methods (Sutton and Barto, 2018)) or the optimality of the action policy itself (policy-based methods (Agarwal et al., 2021)) or the hybrid approaches (Mnih et al., 2016). We next formulate the CAT problem as an MDP. Here are the definitions of the key elements in RL within an adaptive learning/testing system and the overall RL framework of CAT is in Figure 6:
@@
For modeling the complex interactions between examinees and questions during the practical test, Zhuang et al. (2022b) recently design a transformer-based Q-Network, which can identify and extract the perturbations in examinee performance (e.g., guess and slip factors). Stochastic Shortest Path Formulation. Furthermore, CAT can be defined as a Stochastic Shortest Path (SSP) problem (Bertsekas and Tsitsiklis, 1991), which is a special case of MDP. In an SSP, the objective is to find the shortest path (i.e., the minimum test step) from a given initial state $s_{0}$ to goal states. In CAT, the goal state typically represents the completion of the test or the attainment of a predetermined level of proficiency estimation precision. The approach by Gilavert and Freire (2022) uses Linear Programming to find the optimal testing policy $\pi^{*}$, treating the problem like a flow network where each state must have balanced inflow and outflow, except for the start and end points. It denotes variables $x_{s,a}$ as the expected accumulated occurrence frequency for every pair (state $s\in S$, action $q\in\mathcal{Q}$), and equalizes $in(s)$ and $out(s)$ flow model for every state $s$. The flow into a state $s$ is the sum of the expected frequencies of all actions in all other states $s^{\prime}$ that lead to $s$: ${in}(s)=\sum_{s^{\prime},q}x_{s^{\prime},q}P(s|s^{\prime},q)$. The flow out of a state $s$ is the sum of the expected frequencies of all actions in state $s$: ${out}(s)=\sum_{q}x_{s,q}$. The objective function is to maximize the total expected reward $r$, which is the sum of the expected frequencies times the immediate rewards $r(s,q)$ for all state-action pairs: $\min_{x_{s,q}}\sum_{s\in S,q\in\mathcal{Q}}{x_{s,q}r(s,q)}$. Thus the optimal policy $\pi^{*}$ can be obtained by:
@@
Partial-Observable MDP Formulation. Partial-Observable MDP (POMDP) is a generalization of MDP to handle situations where the agent does not fully observe the environment (Hoerger and Kurniawati, 2021). Previously, it assumes that the proficiency of an examinee can be perfectly inferred from their previous responses, then CAT can be modeled as an MDP, where the proficiency estimate can be the states. However, in a more general scenario, the proficiency cannot be perfectly inferred from their previous responses. Because there is always some uncertainty involved (Zhuang et al., 2022a; Drousiotis et al., 2021):
@@ 5.4. Meta Learning Algorithms
Figure 7(left) shows the overall meta learning framework. In the inner-level (Eq.(12)), the question in the support set $\mathcal{D}_{s}^{i}$ for examinee $i$ is sequentially selected by $\pi$, according to his/her previous responses; then binary cross-entropy loss $l(\cdot)$ on $\mathcal{D}_{s}^{i}$ is minimized for estimating the proficiency $\hat{\theta}_{i}$ for the outer-level. In the outer-level (Eq.(11)), the loss of the estimate $\hat{\theta}_{i}$ on the query set $\mathcal{D}_{u}^{i}$ is minimized to learn the selection algorithm $\pi$ and the global parameters $\gamma$. The global parameters $\gamma$ refer to the question characteristics and the examinee ability prior. The algorithm $\pi$ is also model-agnostic. It could be adapted to the given CDM ($f$) automatically by optimizing this problem for efficient selection. Once the question selection algorithm is trained, its parameters do not update during the CAT process and adaptively selects the next item based on previous response behaviors. Based on the above bi-level definition, there have been increasing efforts to improve upon it. Ma et al. (2023) propose a flexible optimization framework Decoupled Learning CAT (DL-CAT). The original BOBCAT obtains the parameters of two modules (i.e., examinee proficiency estimation and question selection algorithm) through coupled inner and outer optimizations, i.e., the result of the outer optimization model is used to measure the quality of the inner. DL-CAT devises a ground-truth construction strategy, and a pairwise loss function, allowing these two models to be trained independently; Feng et al. (2023) introduces a constrained version of BOBCAT to address the question exposure and test overlap issues. Yu et al. (2023) recently introduce the collaborative information of examinees in optimizing this bi-level problem, achieving fast convergence of proficiency estimation. Meta-Learning vs Reinforcement Learning. In CAT, meta-learning can be seen as a higher-level learning process that learns how to adapt a general strategy for question selection to specific examinees based on their responses. Actually, this meta-learning CAT can be reframed as a RL problem. Zhuang et al. (2022b) propose NCAT which transforms the meta-learning problem in CAT into a RL problem. Because the test may stop at any step according to different stopping rules, NCAT simplifies the original objective (Eq(11)) in meta-learning and sums all the test steps to minimize the loss:
@@ 5.5. Subset Selection Algorithms
where $\hat{\theta}^{T}=\arg\min_{\theta}{\sum_{(q,y)\in S}{l(y,f(q,\theta))}}$ is the final proficiency estimate when the test ends $(t=T)$ with the corresponding $T$ responses. In contrast to previous sequential selection methods, it essentially does not require perfect selection at each step, but rather emphasizes the accuracy of the final estimate. From a global perspective, CAT essentially is a Subset Selection problem (Miller, 2002), a fundamental challenge in machine learning and optimization. It revolves around choosing a subset of elements $S$ from a larger set $\mathcal{Q}$ that optimizes a particular objective function $F(S)$ while adhering to specific constraints. However, we cannot directly solve the above optimization problem due to the following main challenge: The true proficiency of the examinee, denoted by $\theta_{0}$, is unknown. It is not available in the dataset, which prevents us from directly optimizing or designing the question selection algorithm. To address this, some researchers have tried to approximate the original objective. Mujtaba and Mahapatra (2021) use the standard error of measurement (SEM) to approach the optimization objective. The SEM provides a measure of confidence in an estimate from a test. They propose using SEM as a new target to measure the accuracy of proficiency estimation: $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|\Rightarrow\min_{|S|=T}\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}},$ where $\mathcal{I}_{i}(\hat{\theta})$ is the information (e.g., Fisher Information) for a question $i$ for an examinee with proficiency $\hat{\theta}$, thus the objective function $F(S)=\sqrt{\frac{1}{\sum_{i\in S}{\mathcal{I}_{i}(\hat{\theta}^{T})}}}$. At each step, it uses multi-objective evolutionary algorithms to obtain the set of Pareto-optimal solutions (Deb, 2011) by maximizing precision and minimizing the number of questions.
@@
where $w(i,j)\triangleq d-\mathrm{max}_{\theta\in\Theta}{\|\nabla l_{i}(\theta)-\nabla l_{j}(\theta)\|}$ is the gradient similarity between question pair $(q_{i},q_{j})$ for this examinee, thus the objective function $F(S)=\sum_{i\in\mathcal{Q}}\mathrm{max}_{j\in S}\;w(i,j)$. The core of BECAT’s subset selection strategy is to find a subset $S$ of size $T$ that maximizes the coverage of $\mathcal{Q}$, quantified by the similarity measure $w(i,j)$ between each question $i\in\mathcal{Q}$ and its closest counterpart $j\in S$. This approach, represented in Figure 8, essentially seeks the most representative questions, aligning with prior selection algorithms but under a new, more rigorous theoretical framework. Given the NP-Hard nature of this optimization, BECAT employs a submodular function approximation. A simple greedy algorithm can finally solve this subset selection problem, with BECAT ensuring that the estimate error remains upper-bounded at each step. The subset selection problem in CAT is a fresh direction with significant potential. This method offers a universal and global framework for question selection, applicable across various complex CDMs that can utilize gradient-based estimations, including neural network models. Crucially, it also provides a theoretical basis for estimation error analysis.
@@
Statistical    Algorithms
Simple implementation and efficient operation
Dependent on IRTs and requires expert knowledge for design
Active Learning
Model-agnostic and flexible
Neglect the nuanced information within CDM parameters
Reinforcement Learning
Automatic generation of selection algorithm; Sequential Decision Making
Incurs additional training costs and potential bias from data-driven selection
Meta Learning
Automatic generation of selection algorithm; Fast Adaptation
Incurs additional training costs and potential bias from data-driven selection
Subset Selection
Strong theoretical guarantees for estimation accuracy
Faces challenges in the initial stages of CAT
@@
Table 1 compares these five categories of selection algorithms in CAT systems, highlighting the generality and interpretability of each category, along with their main advantages and limitations. This overview assists researchers in identifying the most suitable algorithm for their CAT applications, balancing efficiency and complexity (a summary of the performance of different methods is presented in Table 3). It is noteworthy that, despite the superior performance demonstrated by the latest machine learning and deep learning methods, they do not currently supplant traditional statistical approaches. Particularly in testing scenarios that prioritize interpretability or efficiency, statistical methods remain predominant.
@@ 7.2. Fairness
Fairness is a topic of profound societal significance in both education and machine learning research fields, sparking numerous discussions and leading to the development of many fairness-aware learning algorithms (Cleary, 1968; Chai and Wang, 2022; Li and Liu, 2022). As a technology with potential applications in high-stakes testing, fairness in CAT is a paramount concern. In CAT, the bias that leads to fairness issues can be introduced through three components: the Cognitive Diagnosis Model (CDM), the selection algorithm, and the question bank:
@@ 7.4. Search Efficiency
In large-scale educational testing, such as the GRE (Mills and Steffen, 2000), the search efficiency of selection algorithms poses a significant challenge. The algorithms typically evaluate the value of each question in a brute-force manner. Consequently, the time complexity of selecting a question in each testing step is linear to the number of questions in the bank $\mathcal{Q}$, denoted as $O(|\mathcal{Q}|)$. This linear relationship is a bottleneck in intelligent educational systems. Many organizations like GMAT (Rudner, 2009) resort to the manual development of filtering rules by experts, a labor-intensive process that somewhat limits the efficacy of CAT. To address the efficiency issue in question selection, researchers have proposed two kinds of solutions: (1) The application of the Particle Swarm Optimization (PSO) algorithm in IRT has proven effective in mitigating computational complexities (Huang et al., 2009). PSO operates with particles concurrently exploring the search space, where each particle represents a potential solution to the optimization problem of question selection. This parallel exploration allows PSO to cover a larger portion of the search space in each iteration, improving the chances of rapidly identifying optimal questions; (2) Leveraging efficient search methods from the realms of recommendation systems and information retrieval to enhance CAT’s search efficiency, e.g., creating of a tree-based index (Ram and Gray, 2012; Zhu et al., 2018). This approach organizes data in hierarchical tree structures, with balanced trees ensuring logarithmic time complexity for search operations. Recently, Hong et al. (2023) propose a Search-Efficient CAT framework, which employs examinee-aware space partitioning to construct a tree-based index. It reduces the search space and eliminates redundant searches in multi-round testing steps, thereby reducing the search complexity from $O(|\mathcal{Q}|)$ to $O(\log|\mathcal{Q}|)$.
@@
Content Control
Exposure
@@
Unbalanced question usage
Mitigates overexposure;
@@
Test Coverage Algorithmic preferences
Comprehensive assessment;
@@
Fairness Bias in CDMs
Skewed training data;
@@
Algorithmic preferences
Reduces disadvantage for certain groups
(Lord, 2012)
Equating
Different selected questions across examinees
Score comparability;
@@
Robustness Noise Resistance
Random variability;
@@
Modeling Uncertainty
Uncertainty in response
Improves accuracy of proficiency estimation
(Veldkamp and Verschoor, 2019)
@@
Search Efficiency
Large question banks;